# Learning Enhanced Ensemble Filters

### v1

@@ 2.5 EnKF-Inspired Probability Measure-Dependent Analysis Maps
We now propose a specific novel filtering methodology, based on the considerations in the previous subsection, leading to the mean-field formulation of the Measure Neural Mapping enhanced Ensemble Filter (MNMEF). We consider a specific form of $\mathfrak{B}_{\textrm{MNM}}$ inspired by the success of the EnKF and describe it here in the mean-field limit. We generalize (2.11) to take the form
@@
$K_{\theta}^{(1)}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}}\left[\left(\widehat{v}-\bar{v}+\widehat{w}_{\theta}\right)\otimes\left(h(\widehat{v})-\bar{h}+\widehat{z}_{\theta}\right)\right],$ (2.23a)
$K_{\theta}^{(2)}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}}\left[\left(h(\widehat{v})-\bar{h}+\widehat{z}_{\theta}\right)\otimes\left(h(\widehat{v})-\bar{h}+\widehat{z}_{\theta}\right)\right],$ (2.23b)
where $\widehat{w}_{\theta}$ and $\widehat{z}_{\theta}$ are trainable correction terms. We apply a neural operator $\mathfrak{F}(\cdot;\theta):\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{y}}\times\mathcal{P}(\mathbb{R}^{d_{b}}\times\mathbb{R}^{d_{y}})\rightarrow\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{y}}$ to define the correction terms as
@@
Remark 2 (Dependencies of $K_{\theta}$ and $\mathfrak{F}$).
The object $K_{\theta}$ defined in (2.21b) depends on the measure $\rho$ and the true observation $y^{\dagger}$. However, according to the detailed calculation steps (2.22),(2.23),(2.24), the parameterized component in $K_{\theta}$ comes from the neural operator $\mathfrak{F}$, which depends on $\widehat{v},h(\widehat{v}),y^{\dagger}$, and $\rho_{h}$. These seemingly different dependencies are in fact consistent. This is because:
  1. 1.
    Since $\widehat{y}=h(\widehat{v})+\eta$ and $\eta\sim\mathsf{N}(0,\Gamma)$, the joint measure $\rho=\operatorname{Law}\bigl{(}(\widehat{v},\widehat{y})\bigr{)}$ contains information equivalent to $\rho_{h}=\operatorname{Law}\bigl{(}(\widehat{v},h(\widehat{v}))\bigr{)}$ and $\Gamma$. In (2.22), we separate the dependence on $\Gamma$, leaving only $\rho_{h}$.
  2. 2.
    For the input $\widehat{v}$ to $\mathfrak{F}$ given by (2.24), since we compute the expectation over $\widehat{v}\sim\widehat{\pi}$ in (2.23), the dependence on $\widehat{v}$ and $h(\widehat{v})$ is absorbed into $\widehat{\pi}$, a marginal of the joint measure $\rho$ that $K_{\theta}$ depends on.
@@
The mean field map $\mathfrak{B}_{\textrm{MNM}}$ (2.21a) is constrained to be affine in both inputs $\widehat{v}$ and $\widehat{y}$. Indeed,
$\mathfrak{B}_{\textrm{MNM}}(\widehat{v},\widehat{y},y^{\dagger},\rho)=\widehat{v}+K_{\theta}\left(y^{\dagger}-\widehat{y}\right).$
@@ 3 Learning Maps on the Space of Probability Measures
In what follows, we show that the popular transformer architecture may be formulated as a mean field mapping on the space of probability measures; this mapping is mean field because it not only acts on but is also parametrized by probability measures. We call such a neural network architecture the Transformer Measure Neural Mapping (TMNM). This architecture will serve as a parametric family of operators of the form $\mathsf{F}:\mathcal{P}(\mathbb{R}^{d_{u}})\times\mathcal{P}(\mathbb{R}^{d_{w}})\times\Theta\to\mathcal{P}(\mathbb{R}^{d_{v}})$ defined via action on $\mu\in\mathcal{P}(\mathbb{R}^{d_{u}})$ and $\nu\in\mathcal{P}(\mathbb{R}^{d_{w}})$ as
@@
In view of Remark 4 we proceed by describing the TMNM architecture in detail as a mapping between spaces of probability measures. This mean field perspective is more general and justifies the deployment of MNM architectures trained with a finite collection of samples to ensembles of different sizes. Nonetheless, in the development that follows, we provide examples illustrating the finite sample empirical setting such as in Remark 4, which will be the focus of later sections.
@@ 3.1 Transformer Measure Neural Mapping
$\mathsf{F}(\mu,\nu)=\mathsf{S}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\cdots\circ\mathsf{S}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{C}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\nu)\circ\mathsf{S}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\cdots\circ\mathsf{S}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}(\mu).$ (3.3)
@@
$m(u)$ $=\frac{1}{d_{u}}\cdot|u|_{1},$ (3.7)
$\sigma^{2}(u)$ $=\frac{1}{d_{u}}\sum_{k=1}^{d_{u}}\bigl{(}u_{k}-m(u)\bigr{)}^{2},$
@@ 3.2 Attention as a Map on Measures
We present the definition of the attention operator as a map between a product of the spaces of probability measures ${\mathcal{P}}(\mathbb{R}^{d_{u}})$ and ${\mathcal{P}}(\mathbb{R}^{d_{w}})$ into another space of probability measures ${\mathcal{P}}(\mathbb{R}^{d_{v}})$  [^2] In this self-contained subsection we use the notation $d_{v}$ to denote the dimension associated with the output space of probability measures. This is not to be confused with the dimension of the state of the dynamical system (2.1) that is subject of the majority of the remainder of the paper.. If considering the map as acting on measures defined on the unbounded domain $\mathbb{R}^{d_{w}}$ we must make assumptions on the input measures themselves, i.e., tail decay assumptions, so that the attention operator is well-defined. Hence, for ease of exposition we first present attention as an operator $\mathsf{A}:{\mathcal{P}}(\Omega_{u})\times{\mathcal{P}}(\Omega_{w})\to{\mathcal{P}}(\mathbb{R}^{d_{v}})$, where $\Omega_{u},\Omega_{w}$ are bounded open subsets of $\mathbb{R}^{d_{u}},\mathbb{R}^{d_{w}}$ respectively; we will then extend the definition to the general case $\mathsf{A}:{\mathcal{P}}(\mathbb{R}^{d_{u}})\times{\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})\to{\mathcal{P}}(\mathbb{R}^{d_{v}})$, where the subset of measures ${\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})\subset{\mathcal{P}}(\mathbb{R}^{d_{w}})$ will be appropriately defined.
@@
$\mathfrak{A}(u;\nu)=\mathbb{E}_{w\sim p(\cdot;u)}Vw,$ (3.10)
@@
$p(dw;u)=\frac{\exp\bigl{(}\langle Qu,Kw\rangle\bigr{)}\nu(dw)}{\int_{\Omega_{w}}\exp\bigl{(}\langle Qu,Kz\rangle\bigr{)}\nu(dz)}.$ (3.11)
@@
For $\delta,K>0$ we define $\mathcal{P}_{\delta,K}(\mathbb{R}^{d_{u}})\subset\mathcal{P}(\mathbb{R}^{d_{u}})$ as a set of measures satisfying the following exponential tail decay condition:
$\mathcal{P}_{\delta,K}(\mathbb{R}^{d_{u}})=\left\{\mu\in\mathcal{P}(\mathbb{R}^{d_{u}}):\,\,\mu\{\omega\in\Omega:|X(\omega)|\geq t\}\leq 2\exp\bigl{(}-t^{1+\delta}/K\bigr{)}\right\}.$
@@
We define the operator $\mathsf{A}:{\mathcal{P}}(\mathbb{R}^{d_{u}})\times{\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})\to{\mathcal{P}}(\mathbb{R}^{d_{v}})$ via its action on inputs $(\mu,\nu)\in{\mathcal{P}}(\mathbb{R}^{d_{u}})\times{\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})$ as a pushforward of the first input measure $\mu\in{\mathcal{P}}(\mathbb{R}^{d_{u}})$ by a transport map $\mathfrak{A}:\mathbb{R}^{d_{u}}\times{\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})\to\mathbb{R}^{d_{v}}$ which is parametrized by the second input measure $\nu\in{\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})$, hence
@@
The transport map $\mathfrak{A}:\mathbb{R}^{d_{u}}\times{\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})\to\mathbb{R}^{d_{v}}$ takes as input a vector $u\in\mathbb{R}^{d_{u}}$ and a probability measure $\nu\in{\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})$ and involves computation of an expectation under a rescaling of the measure $\nu$ which is parametrized by $u$. Namely for $(u,\nu)\in\mathbb{R}^{d_{u}}\times{\mathcal{P}}_{\delta,1}(\mathbb{R}^{d_{w}})$ it holds that
$\mathfrak{A}(u;\nu)=\mathbb{E}_{w\sim p(\cdot;u)}Vw,$ (3.13)
@@
$p(dw;u)=\frac{\exp\bigl{(}\langle Qu,Kw\rangle\bigr{)}\nu(dw)}{\int_{\mathbb{R}^{d_{w}}}\exp\bigl{(}\langle Qu,Kz\rangle\bigr{)}\nu(dz)}.$ (3.14)
@@
The measure $p(du;s)$ is a well-defined probability measure for $\mu\in\mathcal{P}_{\delta,1}(\mathbb{R}^{d_{u}})$.
@@
$\exp\bigl{(}\langle Qu,Kz\rangle\bigr{)}$ $\leq\exp\bigl{(}\|Q\|\|K\||u||z|\bigr{)}$
$\leq\exp\left(\frac{(\|Q\|\|K\||u|)^{\frac{1+\delta}{\delta}}}{(1+\delta)/\delta}\right)\exp\left(\frac{|z|^{{1+\delta}}}{(1+\delta)}\right).$
@@
where $C(u)$ is a constant depending on $u$ that changes from line to line, where in the third line we applied the substitution $t=\exp\bigl{(}\tau^{1+\delta}/(1+\delta)\bigr{)}$ and used $\nu\in\mathcal{P}_{\delta,1}(\mathbb{R}^{d_{w}})$ for the final inequality. ∎
For implementable algorithms it is useful to consider the finite sample case, where we assume we have access to empirical approximations of the measures $\mu\in\mathcal{P}(\mathbb{R}^{d_{u}})$ and $\nu\in\mathcal{P}_{\delta,1}(\mathbb{R}^{d_{w}})$ via a collection of samples, or ensembles, $\{u(1),\ldots,u(N)\}\subset\mathbb{R}^{d_{u}}$ and $\{w(1),\ldots,w(M)\}\subset\mathbb{R}^{d_{w}}$. The following proposition provides a description of the attention operator when applied to empirical measures. The resulting expressions will be useful more generally to formulate the trainable algorithms we will use for data assimilation.
@@
Assume that, for $\mu\in\mathcal{P}(\mathbb{R}^{d_{u}}),\nu\in\mathcal{P}_{\delta,1}(\mathbb{R}^{d_{w}})$, samples $u(1),\cdots,u(N)$ are drawn i.i.d. from $\mu$ and samples $w(1),\cdots,w(N)\sim\nu$ are drawn i.i.d. from $\nu.$ Then for $\mu^{N}=\frac{1}{N}\sum_{j=1}^{N}\delta_{u(j)}$ and $\nu^{M}=\frac{1}{M}\sum_{j=1}^{M}\delta_{w(j)}$ it holds that

### v2

@@ 2.5 EnKF-Inspired Probability Measure-Dependent Analysis Maps
We now propose a specific novel filtering methodology, which we coin as the measure neural mapping enhanced ensemble filter (MNMEF). In this section, we will discuss the mean-field formulation of this approach. We consider a specific form of $\mathfrak{B}_{\textrm{MNM}}$ inspired by the success of the EnKF and describe it here in the mean-field limit. We generalize (2.11) to take the form
@@
$K_{\theta}^{(1)}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}}\left[\left(\widehat{v}-\overline{v}+\widehat{w}_{\theta}\right)\otimes\left(h(\widehat{v})-\overline{h}+\widehat{z}_{\theta}\right)\right],$ (2.23a)
$K_{\theta}^{(2)}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}}\left[\left(h(\widehat{v})-\overline{h}+\widehat{z}_{\theta}\right)\otimes\left(h(\widehat{v})-\overline{h}+\widehat{z}_{\theta}\right)\right],$ (2.23b)
where $\widehat{w}_{\theta}$ and $\widehat{z}_{\theta}$ are trainable correction terms that depend on the state $\widehat{v}$, observation $h(\widehat{v})$, true observation $y^{\dagger}$, and the joint distribution $\rho_{h}$. We apply a neural operator $\mathfrak{F}(\cdot;\theta):\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{y}}\times\mathcal{P}(\mathbb{R}^{d_{b}}\times\mathbb{R}^{d_{y}})\rightarrow\mathbb{R}^{d_{v}}\times\mathbb{R}^{d_{y}}$ to define the correction terms as
@@
Remark 2 (Factoring Our $\Gamma$).
We note that $K_{\theta}$ (2.21b) depends on the probability measure $\rho$ while $\mathfrak{F}$ (2.24) depends on $\rho_{h}$. The relationship between $\rho$ and $\rho_{h}$ is given by
$(\widehat{v},\widehat{y})\sim\rho\;\Leftrightarrow\;\begin{cases}(\widehat{v},\,h(\widehat{v}))\sim\rho_{h},\\ \eta\sim\mathcal{N}(0,\Gamma),\quad\eta\perp\widehat{v},\\ \widehat{y}=h(\widehat{v})+\eta.\end{cases}$ (2.25)
By factoring out the noise covariance in the term $(K_{\theta}^{(2)}+\Gamma)^{-1}$ in (2.22), we use the explicit knowledge of $\Gamma$ in derivation of $K_{\theta}$. Therefore, $\mathfrak{F}$ uses the input $\rho_{h}$ which does not depend on $\Gamma.$
@@
The mean-field map $\mathfrak{B}_{\textrm{MNM}}$ (2.21a) is constrained to be affine in both inputs $\widehat{v}$ and $\widehat{y}$. Indeed,
$\mathfrak{B}_{\textrm{MNM}}(\widehat{v},\widehat{y};y^{\dagger},\rho)=\widehat{v}+K_{\theta}\left(y^{\dagger}-\widehat{y}\right).$
@@ 3 Learning Maps on the Space of Probability Measures
In what follows, we show that the popular transformer architecture may be formulated as a mean-field mapping on the space of probability measures; this mapping is mean-field because it not only acts on but is also parametrized by probability measures. We call such a neural network architecture the transformer measure neural mapping (TMNM). This architecture will serve as a parametric family of operators of the form $\mathsf{F}:{\mathcal{P}}(\mathbb{R}^{d_{u}})\times{\mathcal{P}}(\mathbb{R}^{d_{w}})\times\Theta\to{\mathcal{P}}(\mathbb{R}^{d_{v}})$ defined via action on $\mu\in{\mathcal{P}}(\mathbb{R}^{d_{u}})$ and $\nu\in{\mathcal{P}}(\mathbb{R}^{d_{w}})$ as
@@
In view of Remark 4 we proceed by describing the TMNM architecture in detail as a mapping between spaces of probability measures. This mean-field perspective is more general and justifies the deployment of MNM architectures trained with a finite collection of samples to ensembles of different sizes. Nonetheless, in the development that follows, we provide examples illustrating the finite sample empirical setting such as in Remark 4, which will be the focus of later sections.
@@ 3.1 Transformer Measure Neural Mapping
$\mathsf{F}(\mu,\nu)=\mathsf{S}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\cdots\circ\mathsf{S}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{C}(\mu,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\cdots\circ\mathsf{S}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\circ\mathsf{S}(\nu).$ (3.3)
@@
$m(u)=\frac{1}{d_{u}}\cdot|u|_{1},\qquad\sigma^{2}(u)=\frac{1}{d_{u}}\sum_{k=1}^{d_{u}}\bigl{(}u_{k}-m(u)\bigr{)}^{2},$ (3.7)
@@ 3.2 Attention as a Map on Measures
We present the definition of the attention operator as a map between a product of the spaces of probability measures ${\mathcal{P}}(\mathbb{R}^{d_{u}})$ and ${\mathcal{P}}(\mathbb{R}^{d_{w}})$ into another space of probability measures ${\mathcal{P}}(\mathbb{R}^{d_{v}})$  [^2] In this self-contained subsection we use the notation $d_{v}$ to denote the dimension associated with the output space of probability measures. This is not to be confused with the dimension of the state of the dynamical system (2.1) that is subject of the majority of the remainder of the paper.. If considering the map as acting on measures defined on the unbounded domain $\mathbb{R}^{d_{w}}$ we must make assumptions on the input measures themselves, i.e., tail decay assumptions, so that the attention operator is well-defined. Hence, for ease of exposition we first present attention as an operator $\mathsf{A}:{\mathcal{P}}(\Omega_{u})\times{\mathcal{P}}(\Omega_{w})\to{\mathcal{P}}(\mathbb{R}^{d_{v}})$, where $\Omega_{u},\Omega_{w}$ are bounded open subsets of $\mathbb{R}^{d_{u}},\mathbb{R}^{d_{w}}$ respectively; we will then extend the definition to the general case $\mathsf{A}:{\mathcal{P}}(\mathbb{R}^{d_{u}})\times{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})\to{\mathcal{P}}(\mathbb{R}^{d_{v}})$, where the subset of measures ${\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})\subset{\mathcal{P}}(\mathbb{R}^{d_{w}})$ consisting of all probability measures whose tails decay strictly faster than exponential, will be appropriately defined.
@@
$\mathfrak{A}(u;\nu)=\mathbb{E}_{w\sim p(\cdot;u,\nu)}Vw,$ (3.10)
@@
$p(dw;u,\nu)=\frac{\exp\bigl{(}\langle Qu,Kw\rangle\bigr{)}\nu(dw)}{\int_{\Omega_{w}}\exp\bigl{(}\langle Qu,Kz\rangle\bigr{)}\nu(dz)}.$ (3.11)
@@
For $\delta>0$ we define $\mathcal{P}_{\delta}(\mathbb{R}^{d_{u}})\subset\mathcal{P}(\mathbb{R}^{d_{u}})$ as a set of measures satisfying the following exponential tail decay condition:
$\mathcal{P}_{\delta}(\mathbb{R}^{d_{u}})=\left\{\mu\in\mathcal{P}(\mathbb{R}^{d_{u}}):\,\,\mu\{\omega\in\Omega:|X(\omega)|\geq t\}\leq 2\exp\bigl{(}-t^{1+\delta}\bigr{)}\right\}.$
@@
We define the operator $\mathsf{A}:{\mathcal{P}}(\mathbb{R}^{d_{u}})\times{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})\to{\mathcal{P}}(\mathbb{R}^{d_{v}})$ via its action on inputs $(\mu,\nu)\in{\mathcal{P}}(\mathbb{R}^{d_{u}})\times{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})$ as a pushforward of the first input measure $\mu\in{\mathcal{P}}(\mathbb{R}^{d_{u}})$ by a transport map $\mathfrak{A}:\mathbb{R}^{d_{u}}\times{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})\to\mathbb{R}^{d_{v}}$ which is parametrized by the second input measure $\nu\in{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})$, hence
@@
The transport map $\mathfrak{A}:\mathbb{R}^{d_{u}}\times{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})\to\mathbb{R}^{d_{v}}$ takes as input a vector $u\in\mathbb{R}^{d_{u}}$ and a probability measure $\nu\in{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})$ and involves computation of an expectation under a rescaling of the measure $\nu$ which is parametrized by $u$. Namely for $(u,\nu)\in\mathbb{R}^{d_{u}}\times{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})$ it holds that
$\mathfrak{A}(u;\nu)=\mathbb{E}_{w\sim p(\cdot;u,\nu)}Vw,$ (3.13)
@@
$p(dw;u,\nu)=\frac{\exp\bigl{(}\langle Qu,Kw\rangle\bigr{)}\nu(dw)}{\int_{\mathbb{R}^{d_{w}}}\exp\bigl{(}\langle Qu,Kz\rangle\bigr{)}\nu(dz)}.$ (3.14)
@@
The measure $p(du;s)$ is a well-defined probability measure for $\nu\in\mathcal{P}_{\delta}(\mathbb{R}^{d_{u}})$.
@@
$\exp\bigl{(}\langle Qu,Kz\rangle\bigr{)}\leq\exp\bigl{(}\|Q\|\|K\||u||z|\bigr{)}\leq\exp\left(\frac{(\|Q\|\|K\||u|)^{\frac{1+\delta}{\delta}}}{(1+\delta)/\delta}\right)\exp\left(\frac{|z|^{{1+\delta}}}{(1+\delta)}\right).$ (3.15)
@@
where $C(u)$ is a constant depending on $u$ that changes from line to line, and where in the third line we applied the substitution $t=\exp\bigl{(}\tau^{1+\delta}/(1+\delta)\bigr{)}$ and used $\nu\in\mathcal{P}_{\delta}(\mathbb{R}^{d_{w}})$ for the final inequality. ∎
For implementable algorithms it is useful to consider the finite sample case, where we assume we have access to empirical approximations of the measures $\mu\in{\mathcal{P}}(\mathbb{R}^{d_{u}})$ and $\nu\in{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})$ via a collection of samples, or ensembles, $\{u(1),\ldots,u(N)\}\subset\mathbb{R}^{d_{u}}$ and $\{w(1),\ldots,w(M)\}\subset\mathbb{R}^{d_{w}}$. The following proposition provides a description of the attention operator when applied to empirical measures. The resulting expressions will be useful more generally to formulate the trainable algorithms we will use for data assimilation.
@@
Assume that, for $\mu\in{\mathcal{P}}(\mathbb{R}^{d_{u}}),\nu\in{\mathcal{P}}_{\delta}(\mathbb{R}^{d_{w}})$, samples $u(1),\cdots,u(N)$ are drawn i.i.d. from $\mu$ and samples $w(1),\cdots,w(M)\sim\nu$ are drawn i.i.d. from $\nu.$ Then for $\mu^{N}=\frac{1}{N}\sum_{j=1}^{N}\delta_{u(j)}$ and $\nu^{M}=\frac{1}{M}\sum_{j=1}^{M}\delta_{w(j)}$ it holds that